OpenStudy (anonymous):

2tv''-v'=0, solve the differential equation. How to solve for v'(t) by seperation of variable?

OpenStudy (anonymous):

help... SOS!!!

OpenStudy (helpmeplease):

I think you have to use reduction of order first.

OpenStudy (anonymous):

Yeah, try setting \[u(t)=\frac{dv}{dt}\rightarrow \frac{d^2v}{dt^2}=\frac{du}{dt}\]Plug in and go to town.